OpenStudy (anonymous):

Write the equation of a line that passes through the points ( -2, 5) and (-2, 11).

OpenStudy (anonymous):

first calculate slope then use the point-slope equation

OpenStudy (anonymous):

change in Y over change of X (11-5)/(-2-(-2)= -3/2 (this is your slope) then use any point you want and plug into point slope formula

OpenStudy (anonymous):

\[y-5= -3/2(x-(-2)\]